OpenStudy (anonymous):

If the radius of a circle is 5 centimeters, how long is the arc subtended by an angle measuring 60°? A) 3/5 π cm B) 5/2 π cm C) 5/3 π cm D) 5/6 π cm

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

sixty degrees is \(\frac{\pi}{3}\) measured in radians

OpenStudy (anonymous):

solve \[\frac{s}{5}=\frac{\pi}{3}\] for \(s\)

OpenStudy (anonymous):

thanks